OpenStudy (anonymous):

3 - 2(x - 4) = 5x + 3

OpenStudy (blurbendy):

3 - 2(x-4) = 5x + 3 3 -2x + 8 = 5x + 3 -2x + 11 = 5x + 3 -7x = -8 x = 8/7

OpenStudy (anonymous):